# A note on Johnson’s rule for minimizing makespan in the Two-Machine Flow Shop scheduling problem

### v1

@@ Abstract
We consider Johnson’s rule for minimizing the makespan in the two-machine flow shop scheduling problem. We show that although the worst-case complexity of Johnson’s rule is $O(n\log n)$, since it requires a complete sorting of the jobs, it is possible to detect in linear time whenever a full sort can be avoided and the optimal solution can be computed in linear time. Computational testing indicates that the linear-time complexity always occurs in practice on standard benchmark instances with uniform distribution of the processing times.
@@ 1 Introduction
In this short note, we consider the classical two-machine flow shop scheduling problem with the objective of minimizing the makespan, commonly denoted by $F2\;||\;C_{\max}$ [3]. A set $J=\{1,\dots,n\}$ of jobs must be processed on two machines $M_{1}$ and $M_{2}$ in this order (flow shop). Job $j$ requires processing time $p_{1,j}$ on $M_{1}$ and $p_{2,j}$ on $M_{2}$. All jobs are available at time 0, each machine can handle at most one job at a time and preemption is not allowed. It is well known that there always exists an optimal schedule which is a permutation schedule and can be obtained by the following Johnson’s rule [4] that first schedules the set $A$ of jobs such that $p_{1,j}\leq p_{2,j}$ in nondecreasing of $p_{1,j}$ and then the set $B$ of jobs such that $p_{1,j}>p_{2,j}$ in nonincreasing of $p_{2,j}$ .
@@
Due to the required sorting, Johnson’s rule runs in $O(n\log n)$ time and $O(n)$ space. As mentioned in [7], schedules obtained by Johnson’s rule ”are by no means the only schedules that are optimal for $F2\;||\;C_{\max}$. The class of optimal schedules appears to be hard to characterize and data dependent.” Here, we make a step in this direction, by showing that a simple procedure requiring linear time allows to check whether or not a full sort can be avoided and, correspondingly, an optimal solution can be computed in linear time. In that case, a huge set of equivalent optimal sequences is also detected.
The rest of this note proceeds as follows. In Section 2 we show how to detect in linear time subsets of jobs that do not need to be sorted in the search of an optimal solution as any permutation of those jobs would also induce an optimal sequence. Correspondingly, it is possible to verify whether the size of the remaining jobs to be sorted is small enough to guarantee an overall linear time complexity. As a counterpart, we also present a family of instances where a full sort is necessary. Then, in Section 3, computational testing indicates that, on standard uniform distributions, for instances with size $n=100,1000,10000,100000$ and $1000000$, an optimal solution was always reached in linear time. Finally, Section 4 concludes the paper with final remarks.
@@
Let assume, without loss of generality, that there are $n=n_{A}+n_{B}$ jobs that are indexed according to Johnson’s sequence where jobs $1,\dots,n_{A}$ belong to set $A$ and jobs $n_{A}+1,\dots,n_{B}$ belong to set $B$. Let also denote by $p^{\max}_{A}$ ($p^{\max}_{B}$) the largest processing time in subset AA ($B$). Then, let $P_{1}=\sum_{i=1}^{n}p_{1,i}$ be the total processing time on machine $M_{1}$. Similarly, let $P_{2}=\sum_{i=1}^{n}p_{2,i}$ be the total processing time on machine $M_{2}$. Correspondingly, let $P_{1A}=\sum_{i=1}^{n_{A}}p_{1,i}$ ($P_{2A}=\sum_{i=1}^{n_{A}}p_{2,i}$) be the total processing time of the jobs in subset $A$ on machine $M_{1}$ ($M_{2}$) and let $P_{1B}=\sum_{i=n_{A}+1}^{n_{B}}p_{1,i}$ ($P_{2B}=\sum_{i=n_{A}+1}^{n_{B}}p_{2,i}$) be the total processing time of the jobs in subset $B$ on machine $M_{1}$ ($M_{2}$). Finally, when jobs are processed on the machines, for each job $i$, let denote by $C_{1,i}$ ($C_{2,i}$) the related completion time on machine $M_{1}$ ($M_{2}$).
@@
If $P_{1A}\leq P_{2A}-p^{\max}_{A}$, let $k_{A}$ denote the smallest index such that
$\sum_{i=1}^{k_{A}}p_{1,i}\leq\sum_{i=1}^{k_{A}}p_{2,i}-p^{\max}_{A}.$
@@
$\sum_{j=1}^{k_{A}}p_{1,j}\leq\sum_{i=1}^{k_{A}}p_{2,j}-p^{\max}_{A},$
for each $i\in\{k+1,\dots,n_{A}\}$, whatever job $i^{*}$ is selected for position $k+1$, the completion time on the first machine of that job respects the inequality
$C_{1,i^{*}}=\sum_{j=1}^{k_{A}}p_{1,j}+p_{1,i^{*}}\leq\sum_{j=1}^{k_{A}}p_{1,j}+p^{\max}_{A}\leq\sum_{j=1}^{k_{A}}p_{2,j}$
namely no idle time is enforced on machine $M_{2}$ for job $i^{*}$. Furthermore, the gap between the completion times on the two machines for the job in positions $\geq k+1$ can only increase with respect to the gap of the job in position $k$, and correspondingly no idle time can ever occur on machine $M_{2}$ for positions $k+1,\dots,n_{A}$. Consequently, the completion time of the last job of set $A$ on the second machine will be the same regardless of the permutation considered for the jobs in positions $k+1,\dots,n_{A}$. ∎
@@
If $P_{2B}\leq P_{1B}-p^{\max}_{B}$, let $k_{B}$ denote the largest index such that
$\sum_{i=k_{B}}^{n_{B}}p_{2,i}\leq\sum_{i=k_{B}}^{N_{B}}p_{1,i}-p^{\max}_{B}.$
Then, for set $B$, an optimal schedule is composed by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order followed by subsequence $k_{B},\dots,n_{B}$ .
@@
An even stronger result in terms of equivalent optimal schedules, can be deduced by the following properties 5, 6 whenever either $P_{1}\leq P_{2}-p^{\max}$, or $P_{2}\leq P_{1}-p^{\max}$ holds.
@@
If $P_{1}\leq P_{2}-p^{\max}$, let $k_{A}$ denote the smallest index such that
$\sum_{i=1}^{k_{A}}p_{1,i}\leq\sum_{i=1}^{k_{A}}p_{2,i}-p^{\max}_{A}.$
@@
The first part of the proof related to set $A$ trivially holds from Property 2 as $p^{\max}_{A}\leq p^{\max}$. For the second part, as
$P_{1}=P_{1A}+P_{1B}\leq P_{2}-p^{\max}=P_{2A}+P_{2B}-p^{\max},$
@@
If $P_{2}\leq P_{1}-p^{\max}$, let $k_{B}$ denote the argest index such that
$\sum_{i=k_{B}}^{n_{B}}p_{2,i}\leq\sum_{i=k_{B}}^{N_{B}}p_{1,i}-p^{\max}_{B}.$
Then, an optimal schedule is composed by scheduling initially set $A$ sequenced in any order, followed by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order, followed by subsequence $k_{B},\dots,n_{B}$ . Also, if $(n-k_{B}+1)\log(n-k_{B}+1)\leq n$, then the optimal solution of problem $F2\;||\;C_{\max}$ can be computed in linear time.
@@
As a counterpart, we show that, in the worst case, a full sort of both subsets $A$ and $B$ may occur. Indeed, consider a $2$-machine flow shop problem with $2n$ jobs. Here, the first $n$ jobs belong to subset $A$ and the processing times are as follows: $\forall j\in 1,\dots,n$, $p_{1,j}=j<p_{2,j}=j+1$. Then, the second $n$ jobs belong to subset $B$ and the processing times are as follows: $\forall j\in n+1,\dots,2n$, $p_{1,j}=2n-j+2>p_{2,j}=2n-j+1$. It is straightforward to show that the unique optimum is given by sequence $1,2,\dots,n-1,n,n+1,\dots 2n-1,2n$ and thus a complete sorting of subset $A$ (in nondecreasing processing time order on the first machine) and subset $B$ (in nonincreasing processing time order on the second first machine) is necessary. In figure 2, the related instance and relevant unique optimal solution for $n=4$ with makespan $25$ is presented.

### v3

# Revisiting Johnson’s rule for minimizing makespan in the Two-Machine Flow Shop scheduling problem

@@ Abstract
We consider Johnson’s rule for minimizing the makespan in the two-machine flow shop scheduling problem. Although the worst-case complexity of Johnson’s rule is $O(n\log n)$, due to the requirement of fully sorting the jobs, we show that it is possible to detect in $O(n)$ time whether a complete sort can be avoided, and an optimal solution can be computed in linear time. For uniformly distributed processing times, a probabilistic analysis shows that linear-time complexity is achieved with probability at least $1-10^{-3}$ for instances of size $n=100$ and at least $1-10^{-9}$ for instances of size $n=200$. Computational experiments further indicate that linear-time complexity is consistently observed in practice on standard benchmark instances with uniformly distributed processing times, as well as on instances generated according to other widely used distributions.
@@ 1 Introduction
We consider the classical two-machine flow shop scheduling problem with the objective of minimizing the makespan, commonly denoted by $F2\;||\;C_{\max}$ [5]. A set $J=\{1,\dots,n\}$ of jobs must be processed on two machines $M_{1}$ and $M_{2}$ in this order (flow shop). Job $j$ requires processing time $p_{1,j}$ on $M_{1}$ and $p_{2,j}$ on $M_{2}$. All jobs are available at time 0, each machine can handle at most one job at a time and preemption is not allowed. It is well known that there always exists an optimal schedule which is a permutation schedule and can be obtained by the following Johnson’s rule [7] that first schedules the set $A$ of jobs such that $p_{1,j}\leq p_{2,j}$ in nondecreasing of $p_{1,j}$ and then the set $B$ of jobs such that $p_{1,j}>p_{2,j}$ in nonincreasing of $p_{2,j}$ .
@@
Due to the required sorting, Johnson’s rule runs in $O(n\log n)$ time and $O(n)$ space. As mentioned in [10], schedules obtained by Johnson’s rule ”are by no means the only schedules that are optimal for $F2\;||\;C_{\max}$. The class of optimal schedules appears to be hard to characterize and data dependent.” To this extent, several papers focusing on the generation of equivalent optimal solutions were proposed. The characterization and enumeration of all optimal sequences was first addressed in [2]. Their work was later extended in [4], where a new sufficient condition of optimality was proposed. More recently, a refinement on the optimality conditions and additional structural insights into the set of optimal schedules were proposed in [6].
Here, we explore general conditions for optimally solving the problem in linear time. More specifically, we prove that a simple procedure requiring linear time allows to check whether or not a full sort can be avoided and, correspondingly, an optimal solution can also be computed in linear time. In that case, in addition, a huge set of equivalent optimal sequences is detected. To the best of our knowledge, this is the first work showing that Johnson’s rule can be implemented in linear time with high probability under standard distributions.
The rest of the paper proceeds as follows. In Section 2 we show how to detect in linear time subsets of jobs that do not need to be sorted in the search of an optimal solution as any permutation of those jobs would also induce an optimal sequence. Correspondingly, it is possible to verify whether the size of the remaining jobs to be sorted is small enough to guarantee an overall linear time complexity. As a counterpart, we also present a family of instances where a full sort is necessary. Section 3 provides a probabilistic analysis on instances with uniform distributions of the processing times showing that linear-time complexity is guaranteed with probability no less than $1-1.5\cdot 10^{-4}$ for instances with size $n\geq 100$ and no less than $1-10^{-9}$ for instances with size $n\geq 200$. In Section 4, computational testing indicates that, on standard uniform distributions, for size $n=100,1000,10000$ and $100000$, an optimal solution was always reached in linear time. Same behavior was observed on other instances with size $n=1000$ for various other widely used distributions. Finally, Section 5 concludes the paper with final remarks.
@@
Let assume, without loss of generality, that there are $n=n_{A}+n_{B}$ jobs that are indexed according to Johnson’s sequence where jobs $1,\dots,n_{A}$ belong to set $A$ and jobs $n_{A}+1,\dots,n$ belong to set $B$. Let denote by $p^{\max}$ the largest processing time in the jobset. Let also denote by $p^{\max}_{A1}$ ($p^{\max}_{B2}$) the largest processing time in subset AA ($B$) on machine $M_{1}$ ($M_{2}$).Then, let $P_{1}=\sum_{i=1}^{n}p_{1,i}$ be the total processing time on machine $M_{1}$. Similarly, let $P_{2}=\sum_{i=1}^{n}p_{2,i}$ be the total processing time on machine $M_{2}$. Correspondingly, let $P_{1A}=\sum_{i=1}^{n_{A}}p_{1,i}$ ($P_{2A}=\sum_{i=1}^{n_{A}}p_{2,i}$) be the total processing time of the jobs in subset $A$ on machine $M_{1}$ ($M_{2}$) and let $P_{1B}=\sum_{i=n_{A}+1}^{n}p_{1,i}$ ($P_{2B}=\sum_{i=n_{A}+1}^{n}p_{2,i}$) be the total processing time of the jobs in subset $B$ on machine $M_{1}$ ($M_{2}$). Finally, when jobs are processed on the machines, for each job $i$, let denote by $C_{1,i}$ ($C_{2,i}$) the related completion time on machine $M_{1}$ ($M_{2}$).
Remark 1.
@@
If $P_{1A}\leq P_{2A}-p^{\max}_{A1}$, let $k_{A}$ denote the smallest index such that
$p_{1,k_{A}}<p_{1,k_{A}+1}\;\;\land\sum_{i\in A:p_{1,i}\leq p_{1,k_{A}}}p_{1,i}\leq\sum_{i\in A:p_{1,i}\leq p_{1,k_{A}}}p_{2,i}-p^{\max}_{A1}.$
@@
$\sum_{j=1}^{k_{A}}p_{1,j}\leq\sum_{i=1}^{k_{A}}p_{2,j}-p^{\max}_{A1},$
for each $i\in\{k_{A}+1,\dots,n_{A}\}$, whatever job $i^{*}$ is selected for position $k_{A}+1$, the completion time on the first machine of that job respects the inequality
$C_{1,i^{*}}=\sum_{j=1}^{k_{A}}p_{1,j}+p_{1,i^{*}}\leq\sum_{j=1}^{k_{A}}p_{1,j}+p^{\max}_{A1}\leq\sum_{j=1}^{k_{A}}p_{2,j}$
namely no idle time is enforced on machine $M_{2}$ for job $i^{*}$. Furthermore, the gap between the completion times on the two machines for the job in positions $\geq k_{A}+1$ can only increase with respect to the gap of the job in position $k_{A}$, and correspondingly no idle time can ever occur on machine $M_{2}$ for positions $k_{A}+1,\dots,n_{A}$. Consequently, the completion time of the last job of set $A$ on the second machine will be the same regardless of the permutation considered for the jobs in positions $k_{A}+1,\dots,n_{A}$. The first condition ($p_{1,K_{A}}<p_{1,k_{A}+1}$) is added to ensure that every job $l\in A$ with $p_{1,l}=p_{1,K_{A}}$ is included in the subset to be sorted.
@@
If $P_{2B}\leq P_{1B}-p^{\max}_{B2}$, let $k_{B}$ denote the largest index such that
$p_{2,k_{B}}>p_{2,k_{B}-1}\;\;\land\sum_{i\in B:p_{2,i}\leq p_{2,k_{B}}}p_{2,i}\leq\sum_{i\in B:p_{2,i}\leq p_{2,k_{B}}}p_{1,i}-p^{\max}_{B2}.$
Then, for set $B$, an optimal schedule is composed by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order followed by subsequence $k_{B},\dots,n$ .
@@
An even stronger result in terms of equivalent optimal schedules, can be deduced by the following properties 5, 6 whenever either $P_{1}\leq P_{2}-p^{\max}_{B2}$, or $P_{2}\leq P_{1}-p^{\max}_{A1}$ holds.
@@
If $P_{1}\leq P_{2}-p_{B2}^{\max}$, let $k_{A}$ denote the smallest index such that
$p_{1,k_{A}}<p_{1,k_{A}+1}\;\;\land\sum_{i\in A:p_{1,i}\leq p_{1,k_{A}}}p_{1,i}\leq\sum_{i\in A:p_{1,i}\leq p_{1,k_{A}}}p_{2,i}-p^{\max}_{A1}.$
@@
The sequencing of set AA trivially holds from Property 2. For the second part, as
$P_{1}=P_{1A}+P_{1B}\leq P_{2}-p_{B2}^{\max}=P_{2A}+P_{2B}-p_{B2}^{\max},$
@@
If $P_{2}\leq P_{1}-p_{A1}^{\max}$, let $k_{B}$ denote the largest index such that
$p_{2,k_{B}}>p_{2,k_{B}-1}\;\;\land\sum_{i\in B:p_{2,i}\leq p_{2,k_{B}}}p_{2,i}\leq\sum_{i\in B:p_{2,i}\leq p_{2,k_{B}}}p_{1,i}-p^{\max}_{B2}.$
Then, an optimal schedule is composed by scheduling initially set $A$ sequenced in any order, followed by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order, followed by subsequence $k_{B},\dots,n$ . Also, if $(n-k_{B}+1)\log(n-k_{B}+1)\leq n$, then the optimal solution of problem $F2\;||\;C_{\max}$ can be computed in linear time.
@@
Remark 2.
We remark that, in Properties 2, 5, there may exist several other jobs $h<k_{A}$ with $p_{1,h}=p_{1,k_{A}}$. Let denote by $k^{\prime}_{A}$ the index of the first of these jobs. Then, to get an optimal schedule, it is sufficient to sort only jobs $1\dots k^{\prime}_{A}-1$ and sequence them in that order, followed by jobs $k^{\prime}_{A}\dots k_{A}$ sequenced in any order (as they have the same processing time on the first machine), followed by jobs $k_{A}+1,\dots,n_{A}$ sequenced in any order. Similar consideration holds for Properties 3, 6, where there may be several other jobs $l>k_{B}$ with $p_{2,l}=p_{2,k_{B}}$ and we denote by $k^{\prime}_{B}$ the index of the last of these jobs. Correspondingly, only jobs $k^{\prime}_{B}+1\dots n$ need to be sorted in subset $B$ preceded by jobs $k_{B}\dots k^{\prime}_{B}$ sequenced in any order preceded by jobs $n_{A}+1,\dots,k_{B}-1$ sequenced in any order.
As a counterpart, we show that, in the worst case, a full sort of both subsets $A$ and $B$ may occur. Indeed, consider a $2$-machine flow shop problem with $2n$ jobs. Here, the first $n$ jobs belong to subset $A$ and the processing times are as follows: $\forall j\in 1,\dots,n$, $p_{1,j}=j<p_{2,j}=j+1$. Then, the second $n$ jobs belong to subset $B$ and the processing times are as follows: $\forall j\in n+1,\dots,2n$, $p_{1,j}=2n-j+2>p_{2,j}=2n-j+1$. It is straightforward to show that the unique optimum is given by sequence $1,2,\dots,n-1,n,n+1,\dots 2n-1,2n$ and thus a complete sorting of subset $A$ (in nondecreasing processing time order on the first machine) and subset $B$ (in nonincreasing processing time order on the second first machine) is necessary. In figure $2$, the related instance and relevant unique optimal solution for $n=4$ with makespan $25$ is presented.
@@
3 Probabilistic analysis
In this section, we perform a probabilistic analysis for instances with $n$ jobs, where processing times are independently and uniformly distributed in the range $\{1,\dots,n\}$. Hence, $p_{\max}=n$. We focus on parameter $k_{A}$ and subset $A$; by the reversibility Property 1, the same analysis applies symmetrically to $k_{B}$ and subset $B$.
Our goal is to analyze the probability of identifying a value of $k_{A}$ sufficiently small to ensure linear complexity in $n$, and to study how this probability behaves as $n$ increases.
3.1 Step 1: Lower bound on the size of set AA
We fix a parameter $\alpha$ with $0<\alpha<1$. For each job $i$, we consider the following restrictive condition:
$C_{1}(i):=\bigl(p_{1,i}\leq\alpha n\bigr)\;\land\;\bigl(p_{2,i}>\alpha n\bigr).$
Clearly, $C_{1}(i)$ implies that job $i$ belongs to set $A$, although the converse does not necessarily hold.
Lemma 1 (Probability of condition $C_{1}$).
For any job $i$, the probability that condition $C_{1}(i)$ holds is
$\Pr[C_{1}(i)]=\alpha(1-\alpha).$
Proof.
Since $p_{1,i}$ and $p_{2,i}$ are independent and uniformly distributed in $\{1,\dots,n\}$, we have
$\Pr(p_{1,i}\leq\alpha n)=\alpha\quad\text{and}\quad\Pr(p_{2,i}>\alpha n)=1-\alpha.$
Let $X$ be the random variable counting the number of jobs satisfying condition $C_{1}$. Then $X$ follows a binomial distribution with parameters $n$ and $p=\alpha(1-\alpha)$. By the Bernoulli binomial formula,
$\Pr(X=k)=\binom{n}{k}[\alpha(1-\alpha)]^{k}[1-\alpha(1-\alpha)]^{n-k}.$
Since we are interested in the probability that at least $k_{A}$ jobs satisfy $C_{1}$, we define
$P_{1}^{*}:=\Pr(X\geq k_{A})=\sum_{k=k_{A}}^{n}\binom{n}{k}[\alpha(1-\alpha)]^{k}[1-\alpha(1-\alpha)]^{n-k}.$
3.2 Step 2: Lower bound on the sum of processing times differences
From Step 1, it follows that with probability at least $P_{1}^{*}$, there exist at least $k_{A}$ jobs $i\in A$ such that $p_{1,i}\leq\alpha n$. Consider the first $k_{A}$ jobs in set $A$ according to Johnson’s sequence. The inequality $p_{1,i}\leq\alpha n$ necessarily holds for all these jobs.
For each such job $i$, define
$\delta_{i}:=p_{2,i}-p_{1,i}.$
Since $p_{1,i}\leq\alpha n$, the variable $\delta_{i}$ is uniformly distributed over
$\{0,\dots,p_{\max}-p_{1,i}\}.$
Let $p^{\prime}:=(1-\alpha)p_{\max}$. We further restrict the support of $\delta_{i}$ to the smaller set $\{0,\dots,p^{\prime}\}$. Under the above restriction, the variables $\delta_{1},\dots,\delta_{k_{A}}$ are independent and uniformly distributed over $\{0,\dots,p^{\prime}\}$ and we can compute a lower bound on the probability that $\sum_{i=1}^{k_{A}}\delta_{i}\geq p_{\max}$ by dynamic programming.
Let $P(j,\gamma)$ denote the probability that $\sum_{i=1}^{j}\delta_{i}=\gamma$, namely, the number of realizations yielding exactly sum $\gamma$ divided by $(p^{\prime}+1)^{j}$.
By considering the restricted distribution, the maximum reachable value of $\gamma$ is $k_{A}p^{\prime}$. The dynamic program is defined as follows:
  • Initial conditions:
    $P(0,0)=1,\qquad P(0,\gamma)=0\quad\forall\gamma>0.$
  • Boundary conditions:
    $P(j,\gamma)=0\quad\text{if }\gamma<0.$
  • Recursion:
    $P(j,\gamma)=\frac{1}{p^{\prime}+1}\sum_{i=0}^{p^{\prime}}P(j-1,\gamma-i),\qquad j=1,\dots,k_{A}.$
Finally, the probability that $\sum_{i=1}^{k_{A}}\delta_{i}\geq p_{\max}$ holds, corresponds to the probability that $\sum_{i=1}^{k_{A}}\delta_{i}<p_{\max}$ fails, which is given by
$P_{2}^{*}=1-\sum_{\gamma=0}^{p_{\max}-1}P(k_{A},\gamma).$
Correspondingly, the product $P^{*}=P_{1}^{*}\cdot P_{2}^{*}$ provides a lower bound on the probability that only the first $k_{A}$ jobs need to be sorted in set $A$. Analogous probability holds for set $B$ and parameter $k_{B}$.
In Table 1, we provide for $n=20n^{\prime}$ with $n^{\prime}=1,\dots 10$ and $p_{\max}=n$, the relevant value of $k_{A}$ with $k_{A}\log k_{A}<n$, the most suitable value of $\alpha$ and the corresponding value of $P^{*}$. The table indicates that $P^{*}>1-10^{-1}$ for $n=40$, $P^{*}>1-10^{-3}$ for $n=100$ and $P^{*}>1-10^{-9}$ for $n=200$, confirming also probabilistically a linear time complexity of the proposed approach when processing times are generated according to a uniform distribution.
$n$ $k_{A}$ $\alpha$ $P^{*}$
2020 44 $0.36$ $0.551$
$40$ 66 $0.41$ $0.912$
$60$ 77 $0.39$ $0.987$
$80$ 99 $0.43$ $1-10^{-3}$
$100$ $10$ $0.42$ $1-10^{-4}$
$120$ $12$ $0.46$ $1-10^{-5}$
$140$ $13$ $0.45$ $1-10^{-6}$
$160$ $14$ $0.44$ $1-10^{-7}$
$180$ $15$ $0.44$ $1-10^{-8}$
$200$ $16$ $0.43$ $1-10^{-9}$
Table 1: Values of $P^{*}$ for increasing value of $n$